[Ian, Daniel]
In spite of my old proclamation, I decided to clean the 4.5 in CF Gate Valve from Alan Rice. All the materials are stainless steel, with two exceptions: some nylon is press fit into stainless steel parts to reduce friction, the gasket to seal the gate valve when closed is made of viton. I did the standard 10 minute sonication in Simple Green:DI Water 1:30, DI water, and isopropanol. I did the following batches: the main body with the CF knife edges, the top with the bellows and screw that allows for actuation, the inside parts (including the parts with nylon, which is fine according to LIGO's methods, section 12.2), and two outside parts which don't touch the vacuum. I am baking at 200°C all purely stainless steel parts that will go into vacuum. On Friday, I will clean the viton according to this procedure, replacing the pressure cook with a shorter sonication and baking out at 120°C for 48 hours. I will finally clean the nylon/steel parts at 80°C for 24 hours.
We need to get cleanish stainless steel screws to assemble the part: 1/4-28 socket head cap screw, 1 5/16” long (qty. 12); 1/4-28 nut (qty. 12); #6-32 socket head cap screw, 3/8” long (qty. 4); #10-32 socket head cap screw, 1/2” long (qty. 1). The 1/4-28 parts can be used from the CF Flange screws (1 1/4" long screws are too short and we have many extra nuts). I got the other screw types from the MCE shop.
A while ago, I contacted Thorlabs technical support on how to make their Polaris parts UHV rated. They generally rate their parts as 1E-5 rated without a bake out and 1E-9 with "proper bake out". They wrote the following and attached the document attached here
For the suggested Polaris bake out procedure, you can refer to the attached document. This will consist of heating the flight hardware in a clean, certified vacuum system (<1 x 10-5 Torr pressure) at the highest temperature permitted without endangering the hardware but at least 10 degree Celsius ( C ) above its in-flight operating extreme, assuming this temperature does not exceed the maximum exposure temperature. Our non-Piezo Polaris mounts have a maximum operating temp at 200° listed on our website, so we can suggest a bakeout at 210°C. The 8-32 and M4 cap screws included with the Polaris mounts are not rated for pressures below 10^-5 Torr. Prior to placing any components in a sensitive vacuum system, a thorough pre-baking in a bake-out oven should be performed to remove all moisture and surface volatiles.
We can't reach 1E-5 Torr in our vacuum oven, but I still think we should bake out the components at ~1 Torr and 200°C.
The pressure seemed stuck at 1.1E-6 Torr, so I vented, took off the 2D MOT, added a spare 2.75" CF blank flange from the Holometer (we have plenty of unopened 2.75" blank flanges) and tightened the screws to ~40 in-lbs. I took off the 12" x 6" plate that was attached to the 2D MOT. I noticed a small hair which I wiped away with an isopropanol wipe but otherwise the inspection looked good. I wiped the vacuum surfaces with an isopropanol wipe, rotated the plate 180°, and installed it, tightening the bolts to 70 in lbs. I turned on the scroll at 2:07 PM and turned on the turbo 6 minutes later when the pressure was 4.7E-2 Torr. The turbo reached full speed 70 seconds later with a power of 4-5 W a few seconds after that. 25 minutes later, the pressure is 8.7E-6 Torr, a bit worse than the past two times but we ultimately care about the final pressure. At 6 PM, the pressure is 2.2E-6 Torr.
This is a first iteration of a power recycled finesse model of the GQuEST demonstrator.
For this initial version, I have the curved mirror as the input mirror. The ROC is ~20m, which yields a beam size on the end mirror of 2mm. The total length of the power recycling cavity is 4.2m (PRCM1->PRCM2: 1.2m,PRCM2->PRCM3: 1.2m,PRCM3->BS: 1.2m, ArmLength:.6m). A propagate beam call of this system shows the input beam size for this configuration would be ~2.3mm. port_tree.svg shows a map of all the input and output ports of the mirrors and beam splitters (this is less useful but I had been unable to get one of these to work before and thought it was interesting).
Next, after talking with Sander, I want to find a range of PRM lengths, IFO lengths, and PRM ROCs that yield a beam waist of 2mm on the end mirror. To simplify this idea and to reduce the calculation time, I have merged PRCM1->PRCM2: 1.2m,PRCM2->PRCM3: 1.2m,PRCM3->BS: 1.2m into one length. Then I calculate if my model is stable over a range of ROC values, IFO lengths, and PRM lengths, combining the two lengths into a single power recycling cavity length. PRCgeometrywithcontours.svg is the result. I've included contour lines to see lines of constant beam size on the end mirrors.
I plan on iterating on this but this was a first attempt. For now I have a working model of the single simplified system and a version of the code we can iterate over specific paramaters. Note that for my simplified version an ROC for the PRM of 1.6m (which is what we have in the lab already) is not stable (also 3m is not stable, the size of the existing spares set aside for the LFC, needs to be greater than about 4.5m).
Since the pressure only reached 1.0E-6 after 2 full days of pumping, I decided to leak check again. There was no leak 180° from the past leak on the plate I rotated, indicating the plate was fine (a visual inspection showed it to be fine as well). There was a leak in the same place as last time (bottom right (south) of the west 6" x 12" plate that goes to the 2D MOT), but maybe smaller, around 2.5E-10 Torr L/s. Interestingly, closing the valve at the end of the leak check spiked the measured leak rate to 4.8E-10 Torr L/s.
I vented the chamber and removed the plate to inspect it and the frame. There was very slight discoloration on the plate in a few places, but I couldn't feel anything. The chamber looked fine. I replaced the plate. Putting in the screws was harder with the 2D MOT in the way, but I don't see an easy way to redesign. We shouldn't have to take this plate to add or adjust components in the vacuum chamber. Due to the awkward angle, I very lightly stripped the top of the head of a screw. This is cosmetic as I could easily get a hex key in, but the screw should be thrown out if we take off the plate again.
I started the scroll at 10:55 PM. Since I suspect the 2D MOT is tugging on the plate through the bellows, I pushed the plates that hold the 2D MOT foward after the pressure was ~1 Torr. I turned off safe start and turned on the turbo at 11:01 PM, the turbo quickly ramped up to speed, maxing out at 99 W drawn. As of writing this at 11:39 PM, the pressure is 3.9E-6 Torr and slowly dropping. If the pressure doesn't get close to 1E-7, I want to take off the 2D MOT, rotate the plate, and add a 2.75" blank flange. We should also use the RGA to confirm the presssure is from a leak and not outgassing. If the leak stays in the same place, the next step would be taking apart the frames, which would be a total disassembly. Being able to valve off the turbo would also make leak checking a lot better, so hopefully the gate valve arrives soon.
Time | Pressure (1E-6 Torr) |
1:01 Monday | 2.2E-6 |
10:20 AM | 1.3E-6 |
3:19 PM | 1.2E-6 |
10:30 AM Tuesday | 1.1E-6 |
Based on this leak checking, there seems to be a singular problematic area. The easiest idea is to rotate the square plate next to the area to see if that fixes anything or the leak moves. There could also be an installation error this would fix.
I turned off the turbo pump, waited for it to stop spinning, then turned off the scroll pump. I turned on the N2 line and slowly opened the needle valve. I waited for the pressure gauge to read over pressure, removed the screws (a bit of air rushed in), rotated the plate 180°, put it back on, tightened at 10, 20, 30, 40, and 70 in-lb (twice). I then turned off needle valve and N2 line. I turned on the scroll at 4:37 PM and the pressure was 6.7E-2 just 5 minutes after turning on scroll, faster than before. I turned on turbo at this point, keeping soft start on for a pure comparison, but we should turn it off soon.
After 16 total minutes, the pressure was 5.5E-6 Torr (slightly better than LFC, ~20x better than Dewer at this time), 5W drawn (LFC 10W, Dewer 16W at this time), 1167 Hz pump speed. I am hopeful this was the fix as we've reached a better pressure much sooner. There is a chance the pumps run better the second time, but it took 2 hours of pumping to reach this presssure last time on the RbQ vacuum chamber.
Time | Pressure (1E-6 Torr) |
5:10 PM | 3.1E-6 |
10:08 AM | 1.3E-6 |
7 PM Saturday | 1.1E-6 |
9:58 PM Sunday | 1.0E-6 |
I made slots on both ends of three 1/4-20 to 3/8-18 thread adapters for use with a flat-head screw driver. I used the band saw meant for steel and held the adapters with a collet (a 15/64" collet and 11/32" collet I think) and a square collet holder. Ian checked that a flat-head screwdriver fits into the slot.
I leak tested the RbQ Vacuum Chamber with LIGO's leak checker. Since the 4.5" gate valve hasn't arrived yet, we had to keep the turbo on which makes leak checking more difficult. However, I did fine at least one leak: the bottom of the rectangular prism to cube interface on the south side (roughly 5 Torr*L/s, but this is pretty arbitrary since the turbo is on (and helium spray level ambiguous)). I also discovered all three screws that secure the windows were loose. I gently tightened them. I turned off the leak checker and the helium line.
I plan on letting the turbo pump a bit longer and seeing if the pressure gets lower. I want to take off, inspect, and reattach these plates since it should be easy. i could rotate the 6" square plate to try to see if that changes the leak spot. The leak could also be caused from the rectangular prism to cube interface itself. We would have to take off 8 plates to fix that, so I want to try that last. If we do that, I think I want to take off the cube entirely and evaluate the pressure with just the rectangular prism. We can make a MOT without the cube, but not a conveyer belt. We could attach the tee and include the ion pump as well as the electronics flange.
[Jeff, Torrey]
After the setup of new monitors in the control room, the USB connection to the Teraxion laser was disrupted. To restore the connection required rebooting the laser while it was USB connected to the computer.
The laser was able to connect via USB to the lab laptop without a reboot, so it seems to be a quirk of our extender that it may require a reboot of the device to properly forward it to the host computer.
The Agilent controller worked fine after this fix. I had to unplug and plug back in the USB to the SRS RGA but now it works as well.
[Ian, Torrey]
We tested how much the lab goggles are blocking light at 775 and 1550 nm. We did this by setting the power meter on the table in an existing path, recording the value, and then putting the goggles just before the power meter and recording the values. We also cleaned the goggles before testing. We did this with 5 different goggles randomly chosen from the box of goggles. For the "dark count" of the, we turned off the light and blocked the beam up stream. I think there was a large amount of scatter because the dark count was always higher than the goggled blocked measurement. The values were as follows:
Wavelength | Input Power | Output Power | OD |
775 nm | 1.75 mW | 6 nW | ~5.5 |
1550 nm | 14 mW | 3 nW | ~6.66 |
The goggles claim to be OD 7+ @ 696-1550nm.
If you saw a previous version of this where we were claiming different numbers, we didn't account for the ambient room light being at uW levels. Turning out the room lights accounts for this.
I grabbed the LIGO Lab leak checker from Maty Lesovsky and put it in B111A to leak check the RbQ Vacuum System. We can also use it to leak check the SNSPD Dewer. Maty said she shouldn't need it for a few weeks.
I added a 2.75" CF to KF25 adapter to the RbQ vacuum chamber composite angle valve in preparation for using the leak checker (which uses a KF25 hose). I tightened the bolts until there was flange to flange contact all around. The pressure didn't change from 2.6E-6 Torr during any of this work.
I plugged in RbQ Vacuum Chamber turbo controller into the wall outlet. After warming up, the pressure was "over pressure". I enabled soft start, hit show decrease speed, changed the gas load to air (I don't think this does anything), and changed the pressure units to Torr. I plugged in scroll into the wall outlet and turned it on at 11:57 AM. It starting pumping after 35 seconds, and the pressure gauge showed numerical readings a few seconds after that. I put the interlock bypass D-Sub into the turbo controller and started turbo at 8E-2 Torr, 6.5 minutes after turning on scroll. While in the soft start, the turbo pump drew up to 95W, around 4E-5 Torr 11.5 total minutes in and 2E-5 13.5 total minutes in. After 18 total minutes, the pressure was 1.6E-5 Torr (2x worse than LFC, 10x better than Dewer, all 12 minutes after starting the turbo), 16 W drawn (LFC 10W, Dewer 16W, all 12 minutes after starting the turbo), 1167 Hz pump speed (this is manually set). The pressure was 7.2E-6 at 1:11 PM. I think we can get below 1E-6 Torr, but this is still too high for an Rb MOT, which should be near 1E-8 Torr (the Berkeley class lab operates aaround 4E-8 Torr). The cube baseline pressure is 1E-7 Torr. We probably need to leak check.
Time | Pressure (1E-6 Torr) |
---|---|
1:59 | 5.5 |
2:45 | 4.7 |
3:26 | 4.1 |
3:54 | 3.8 |
4:17 | 3.9 |
4:26 | 3.8 |
10:48 | 2.6 |
By eye, I think we will get to 1-2E-6 Torr. I will get the leak checker tomorrow and hope to use it soon.
[Ian, Torrey]
We added the two-monitor setup to the control room. This should allow us to have more real estate to see more screens in the Moku software. We also switched the computers' positions, so now Gouy is on the left. This is because Gouy is the only one that has two HDMI ports. I also tidied up some of the cables that were in the cable tray.
I have noted the following voltages on the Laser Filter Cavity (LFC) Ion Pump Controller and the pressure according to the Agilent FRG702 pressure gauge. The voltage corresponds to the current draw by the ion pump. These are only from "steady" times, not when there is a sudden change in the system, for example turning the ion pump on or closing the gate valve to the turbo pump. I graphed this data. The fit appears linear, but the lowest voltage (not steady state) I've seen is 2.1 V when first turning on the controller, so the fit seems good up to ~1.4E-6 Torr. Maybe this is when the current maxes out, assuming a constant current to voltage conversion.
Pressure (Torr) | Voltage (V) |
4.7E-07 | 3.95 |
5.1E-07 | 3.93 |
4.4E-07 | 4.09 |
3.1E-07 | 4.29 |
3.0E-07 | 4.32 |
2.5E-07 | 4.44 |
2.2E-07 | 4.52 |
1.4E-07 | 4.67 |
2.4E-07 | 4.36 |
3.0E-07 | 4.22 |
Updated data. It looks like there are two linear regions. I would expect a linear low pressure region and some rolloff at higher pressures as the ion pump satruates, but I expected the rolloff to be above 1E-6 Torr because the ion pump works at and under 1E-6 Torr.
This is a continuation of LeeLog 12105. TLDR - Teraxion laser is fully installed and functioning correctly except for the modulation port required to lock a cavity.
We put the swap on hold as a result of there being no internal Faraday Isolator in the teraxion seeder. Lee mentioned this should be fine for the short term. We plan on buying an isolator soon. In the mean time we should be fine to run this test.
Notes for the swap:
-Powered down amplifier. Turned off Thorlabs seeder.
-Secured laser module to the breadboard.
-Connected the module via USB extension to the BREWSTER computer. Connected fiber output to the power meter via fiber adapter. Turned power on.
-Installed the teraxion LXM laser control GUI on the BREWSTER computer.
-We successfully turned on the laser via the LXM laser control software. Note the software has an interlock system with a password, this password is 1234 and can be found in the manual. The initial output of the laser is 45mW as seen in the photo. I have concerns with controlling the laser via this little GUI. I don't know how closing the program or disconnecting the USB effects the continuous operation of the laser. This is important to ensure that no damage is done to the amplifier. We may test how easy it is to disrupt operation of the laser module before connecting to the amplifier. Testing in the software, there is no "are you sure button" when disabling the output. One click disables it.
-The laser has ADCs and a popout in the software to track parameters of the laser like TEC current, laser temp, etc. See adc.png.
-I decided to test the continuous operation of the laser. Click the "Exit" button in the bottom right of the software does not disrupt lasing. I will assume that the laser will not shut down unless commands in the software are giving. This alleviates my worry outlined above.
-I plan on replacing the fiber PBS with a 75:25 fiber BS. This means one laser will have significantly reduced output. I need to double check the nominal output of the thorlabs laser to see if output*.25 is high enough to properly seed the amplifier.
-Replaced the fiber PBS with the 75:25. The output for the thorlabs seeder after the :25 path in the fiber beam splitter is ~10 mW (using the power meter, this value will change for the amplifier threshold power detector), which is most likely below the required input for the amplifier. Reminder: "Turn the Key ON: if the input power is lower than the pre-determined lower threshold power (typically 10-20 mW), there is an Alarm LIP (low Input Power), otherwise the preamplifier turns ON." Alternatively, we can swap the outputs of the fiber BS to avoid any alarms. The white output fiber should be used for the teraxion, the red output fiber for the thorlabs. Otherwise just getting one of these but a 50:50 would avoid having to ever swap any fiber connections. This is recommended.
-The teraxion is now seeding the amplifier. The power meter read 30.3 mW, the amplifier seeder is reading 27 mW. Both of the values are above the threshold. Continuing with turning on Amplifier.
-Turned key and hit enable on amplifier. The amplifier is now in pre-amplification mode. ~300 mW at the output, seeded by the teraxion.
-There is no 775 light. No light seems to be exiting the SHG. I suspect the wavelengths of the 2 lasers are different enough that the temperature controller needs to be tuned to the proper temperature for the SHG. I don't need 775 light for this test however so I will just note it and move on. Alternatively the wavelength of the laser can be tuned via a temperature controller. You most likely can tune to the same wavelength as the thorlabs controller so the SHG temperature controller doesn't have to be touched.
-The last step is we need to be able to modulate on the new laser. The manual says "please refer to the device test sheet to determine if your laser module accept ±2.5 V input or 0-4 V". Our test sheet on the wiki says +/- 2.5V. If using the moku this will never be a problem as we can't operate above +/- 2V in MIM anyways.
-Connected a moku output to the modulation port of the teraxion laser via patch panel and SMA adapter.
-Everything is connected. Scanning 2 V pp on the modulation port doesn't flash the cavity. I don't think anything is actually scanning. The scan amplitude is .2 GHz. The FSR is 125 MHz. Even at half the scan range we should see something. I tuned the temperature controller to have the cavity naturally flashing, and then turn on the modulation scan and still don't see anything. Manual shows a diagram connected to a waveform generator with a "10-100kHz" signal. Changed the scan frequency to this range, no change. I confirmed the moku is outputting a voltage via T-ing off the output and checking on a different scope. There is a triangle wave.
-Not sure what else to try so I checked the Ultra Narrow Linewidth Mode that is unique to the LXM-U that we have. It works.
Software download info can be found at https://wiki.mccullerlab.com/Main/MX-S24007